# Nonlinearity-induced symmetry breaking in a system of two parametrically driven Kerr-Duffing oscillators

### v1

@@ Abstract
We study the classical dynamics of a system of a pair of Kerr-Duffing nonlinear oscillators, coupled by a nonlinear interaction and subject to a parametric drive. Within a rotating wave approximation (RWA), we analyze the steady-state solutions for the oscillation amplitude of the two oscillators. In the most relevant case of identical oscillators, we separately investigate configurations in which only one oscillator is parametrically driven or both of them are simultaneously driven. In the latter regime, a special attention is paid to the symmetric case where the parametric drives acting on the two oscillators are equal: for an increasing value of the detuning of the parametric drive, a transition to a multi-stable, symmetry-breaking regime is found, where the two oscillators display different oscillation amplitudes and phases.
@@ I Introduction
Nonlinear dynamics of resonantly driven oscillator modes is a very general theme common to many fields of physics, from nanomechanical systems to photonics. In the mechanical context, it has received growing attention in view of metrological applications [1, 2, 3, 4, 5, 6, 7] as well as from a fundamental science perspective. Driven nonlinear nanomechanical systems constitute, in fact, an ideal platform for the investigation of fundamental aspects of nonlinear dynamics and nonequilibrium fluctuations [8]. In photonics, models involving a few resonantly-driven modes coupled by nonlinear processes provide a powerful framework to describe nonlinear optical processes leading to the mixing of different beams and to the generation of new harmonic and/or sub-harmonic components and are at the heart of devices for quantum and/or nonlinear optics applications [9, 10, 11, 12, 13] .
@@
Independently of the specific platform used, nonlinear coupling between linear modes is expected to become strong when the ratio between their resonance frequencies is an integer (linear resonance 1:n) or rational (parametric resonance 2:n), viz. the so-called internal resonances [14]. When internal resonance are not present, nonlinear interaction arises from the dispersive interaction between the modes. If one of the coupled modes is driven in linear regime nonlinear interactions typically manifest as a dispersive shift of the mode frequencies, which depends on the oscillation amplitude of the driven mode. Such nonlinear interaction effects have been observed in a variety of nanomechanical systems, e.g. doubly clamped beams or strings, clamped nano- or microcantilever,s and in doubly clamped, nanomechanical silicon nitride string resonators [15, 16, 17, 18, 19, 20, 21, 22, 23].
@@
In photonics, nonlinear interaction processes are mediated by the nonlinear response of the material medium, in particular second-order $\chi^{(2)}$ and third-order $\chi^{(3)}$ nonlinearities, leading to a variety of phenomena, from optical bistability to parametric amplification and oscillation [9, 11, 10, 12, 13, 24]. In the last decades, growing attention has been paid to systems featuring a complex interplay of several nonlinear phenomena at once and/or well-defined discrete modes [25, 26, 27]. Nowadays, an active frontier is to explore fundamental aspects of non-equilibrium fluctuations around the steady-state of spatially extended systems and shine a light on their relation to fundamental questions of pattern formation in nonlinear systems [28, 29, 30, 31].
@@
In this paper, we investigate a simple model based on two Kerr/Duffing oscillators coupled through nonlinear interaction terms, proportional to the product of the squares of the two amplitudes of the two modes $\sim q_{1}^{2}q_{2}^{2}$. Within the rotating wave approximation (RWA), we analyze the possible stationary, paired solutions and their dynamical stability in the two cases when either only one of the two modes is driven or when both resonators are driven by parametric forces modulating the oscillator frequencies at a frequency close to twice their bare frequency. When both resonators are parametrically driven at the same frequency, the paired amplitudes of the two modes show a complex and intriguing behavior which is also characterized by the presence of several instability ranges in the frequency detuning. As a most interesting result, we find that when the two resonators have equal parameters and are driven with equal amplitudes, the overall symmetry of the system can be spontaneously broken for moderately weak nonlinear interaction, leading to an intriguing multistability phenomenon featuring steady states with different amplitudes and phases on the two oscillators.
@@
The article is organized as follows. In Sec. II, we introduce the model and the effective dynamical equation in the rotating frame using the RWA. In Sec. III, we shortly discuss the behavior of the system when both resonators have non-zero oscillation amplitude in presence when only resonator is driven. Finally IV, we analyze the intriguing case when both resonators are simultaneously driven and the emergence of symmetry breaking.
@@ II The system
We introduce the mathematical model that describes two nonlinear resonant modes coupled by the cross-Kerr / parametric interaction which are driven parametrically. We consider two resonators with equal damping coefficients and equal self-Kerr-Duffing parameters. We refer to the Appendix A for the details.
In the rotating frame and using the RWA, the equations for the complex amplitudes associated with the two modes oscillating at frequencies $\omega_{d}$ are
@@
Here $\delta\Omega_{n}$ corresponds to the scaled detuning for the resonator $n$ and $\delta\Omega_{T,n}$ is proportional to the strength of the parametric drive. It also corresponds to the (scaled) frequency threshold for the self-sustained oscillatory motion of the resonator $n$ in the limit of vanishing damping and without interaction The parameter $g$ represents the scaled coupling strength for the interaction.
@@
Notice that if we have the trivial solution for one of the two resonators $\bar{z}_{n}=0$, the equations for $\delta z_{1}$ and $\delta z_{2}$ become decoupled.
@@
Hereafter we focus our analysis on the strong nonlinear regime in which the motion is dominated by the nonlinear interaction such that we can neglect the damping force, which is valid in the limit $3\kappa A_{n}^{2}/(8\omega_{d})\gg\Gamma$ with $A_{n}$ the amplitude of the oscillatory motion of the resonator $q_{n}(t)=A_{n}\cos(\omega_{d}t+\theta_{n})$, $\kappa$ is the Kerr-Duffiing parameter, $\Gamma$ is the damping and $\omega_{d}$ is half of the drive frequency (see appendix A). This approximation is not valid when we are close to the parametric threshold when the resonator starts to oscillate at frequency $\omega_{d}$ and the amplitude can be small enough to violate the previous condition. However, the effect of the damping is the renormalization of the threshold. This approximation captures qualitatively the behavior of the nonlinear system, although it cannot be quantitatively correct in correspondence with the thresholds.
@@
First of all, for a single parametrically driven nonlinear resonator (uncoupled with $g=0$), one has the following (stable) steady-state solution for $\bar{z}$: the resonator starts to oscillate at frequency $\omega_{d}$ above the detuning threshold $\delta\Omega>-\delta\Omega_{T,1}$ and it is entirely out of phase with respect to the drive, in the limit of vanishing damping (see appendix B). The trivial solution $\bar{z}^{(0)}_{1}=0$ is unstable in the detuning range $|\delta\Omega|<\delta\Omega_{T,1}$.
When we switch on the interaction, we discuss the stationary pair solutions. In particular, the trivial solution $\left(\bar{z}_{1}=0,\bar{z}_{2}=0\right)$ is still a stationary solution and unstable for $|\delta\Omega_{i}|<\delta\Omega_{T,i}$ as for the case $g=0$. On the other hand, the pair solutions $\left(\bar{z}_{1}\neq 0,\bar{z}_{2}=0\right)$ have decoupled equations of the fluctuations Eq. (4) (the coefficients $c$ and $d$ vanish) but they are correlated as the finite value of the amplitude of the driven mode $\bar{z}_{1}$ affects the fluctuation of the second non-driven mode $\delta z_{2}$ (for example the coefficient $a_{2}$ depends on $\bar{z}_{1}$) and therefore the range of stability can be different from the noninteracting case, see appendix C.
@@ III Finite amplitude solutions for one driven resonator
We start the analysis by considering that the external drive is applied only to the first resonator $(\delta\Omega_{T,2}=0)$.
In appendix D, we give a detailed analysis of this regime, in which we report the analytical solutions in table 1. The different possible solutions are associated with different ranges of the coupling strength $g$, in particular in the ranges $30<g<1/$ 3, $11/3<g<$ 1 and $g>1$. We checked that the numerical results are in good agreement with the analytical solutions shown in appendix D.
In short, when only the first mode is driven, the parametric interaction changes the system’s stability range even when we have the trivial solution for the second non-driven mode. Furthermore, the first resonator can drive the state of the second resonator in an oscillatory motion.
We give an example of results for this regime in Fig. 1 (numerical results) in which we plot stable pair solutions (solid lines) and unstable ones (dashed lines). In particular, we notice that, for the stable solutions, the second non-driven resonator starts to oscillate for a given threshold detuning. This leads to a discontinuity in the amplitude curve of the driven resonators. As another possible case (not shown) the curve for the second resonator can have a discontinuity in which the amplitude $\bar{z}_{2}$ has a jump from zero to a finite value.
Figure 1: Example of the behavior for a single-driven mode. Legend: The amplitudes of the two different resonators are distinguished by the symbol, i.e., $\blacksquare$ for $|\bar{z}_{1}|$,
$\blacksquare$
for $|\bar{z}_{2}|$. Lines with the same color show solution pairs of the two resonators. Solid lines correspond to stable solutions. Dashed lines correspond to unstable solutions. Parameters: $g=1/2$, $\delta\Omega_{T,1}=0$ and $\delta\Omega_{T,2}=5$. The blue lines show the solution pairs with only one oscillating resonator while the other one is at rest.
@@ IV.1 Numerical analysis for asymmetric drive $\delta\Omega_{T,1}\neq\delta\Omega_{T,2}$
Figure 2: Example of the behavior of the system for two driven modes with different drive strengths. Legend: The amplitudes of the two different resonators are distinguished by the symbol, i.e. $\blacksquare$ for $|\bar{z}_{1}|$,
$\blacksquare$
for $|\bar{z}_{2}|$. Lines with the same color show solution pairs of the two resonators. Solid lines correspond to stable solutions. Dashed lines correspond to unstable solutions. The green and blue lines are the paired solutions with $\bar{z}_{2}=0$ or $\bar{z}_{1}=0$. Parameters: (a): $\delta\Omega_{T,1}=3$, $\delta\Omega_{T,2}=5$, $g=1/8$. (b): $\delta\Omega_{T,1}=10/3$, $\delta\Omega_{T,2}=5$, $g=1/2$.
On the other hand, for $\delta\Omega_{T,1}\gg\delta\Omega_{T,2}$, for moderate force acting on the first resonator $\delta\Omega_{T,1}\gg 1$ and not so small coupling $g$, the solutions are connected qualitatively to ones obtained for the case of a single-driven mode, see Sec. III. One can estimate this regime in the small damping limit in which we have ${|z_{1}|}^{2}\approx\delta\Omega_{T,1}$ at small detuning. Hence the effects of the first resonator dominate on the second drive for $g\Omega_{T,1}\gg\Omega_{T,2}$. In such regime the interaction between the two resonators dominates over the second drive, the amplitude of the second - weakly driven - resonator is determined by its parametric interaction with the first resonator in a way qualitatively similar to the results discussed in Sec. III.
@@
For $g=1/8$ , we have stable solutions of finite amplitudes, $\bar{z}_{1}\neq 0$ and $\bar{z}_{2}\neq 0$, (turquoise lines) up to some detuning, they disappear, and two new stable solutions emerge again at larger detuning (purple lines). The system has multistability when we regard the phase since the modulus of the solutions shown in Fig. 2 corresponds to solutions of different phases. One can show that, in total, there are four different paired solutions corresponding to the four possible combinations of the phases.
@@
IV.2 Numerical analysis in the symmetric case $\delta\Omega_{T,1}=\delta\Omega_{T,2}$
@@
Figure 3: Example of the behavior of the system for two driven modes with symmetric drive. The orange indicates that two solutions overlap/exist with the same amplitude. The solid lines are the stable solutions, whereas the dashed lines correspond to the unstable ones. Parameters: $\delta\Omega_{T,1}=\delta\Omega_{T,2}=10$ (a): $g=1/5$ (b): $g=3/10$ (c): $g=1/2$.
@@
V Broken-symmetry solutions of the symmetric case

### v2

@@ Abstract
We study the classical dynamics of a system comprising a pair of Kerr-Duffing nonlinear oscillators, which are coupled through a nonlinear interaction and subjected to a parametric drive. Using the rotating wave approximation (RWA), we analyze the steady-state solutions for the amplitudes of the two oscillators. For the case of almost identical oscillators, we investigate separately the cases in which only one oscillator is parametrically driven and in which both oscillators are simultaneously driven. In the latter regime, we demonstrate that even when the parametric drives acting on the two oscillators are identical, the system can transition from a stable symmetric solution to a broken-symmetry solution as the detuning is varied.
@@ I Introduction
Nonlinear dynamics of resonantly driven oscillator modes is a very general theme common to many fields of physics, from nanomechanical systems to photonics. In the mechanical context, it has received growing attention in view of metrological applications [1, 2, 3, 4, 5, 6, 7] as well as from a fundamental science perspective. Driven nonlinear nanomechanical systems constitute, in fact, an ideal platform for the investigation of fundamental aspects of nonlinear dynamics and nonequilibrium fluctuations [8, 9, 10, 11]. In photonics, models involving a few resonantly-driven modes coupled by nonlinear processes provide a powerful framework to describe nonlinear optical processes leading to the mixing of different beams and to the generation of new harmonic and/or sub-harmonic components and are at the heart of devices for quantum and/or nonlinear optics applications [12, 13, 14, 15, 16] .
@@
In general, nonlinear coupling between resonant modes is expected to become strong when the ratio between their resonance frequencies is an integer (linear resonance 1:n) or rational (parametric resonance 2:n), viz. the so-called internal resonances [17]. In particular, an interaction of the kind 2:2, also known as cross-Kerr interaction can be described as the oscillations of the first resonator acting as a parametric frequency modulation on the second resonator. When internal resonances are not present, the 2:2 nonlinear interaction can still give a dispersive shift of the mode frequencies, which depends on the oscillation amplitude of the driven mode. Such a dispersive interaction has been observed in a variety of nanomechanical systems, e.g. doubly clamped beams or strings, clamped nano- or microcantilevers, and in doubly clamped, nanomechanical silicon nitride string resonators [18, 19, 20, 21, 22, 23, 24, 25, 26].
@@
In photonics, nonlinear interaction processes are mediated by the nonlinear response of the material medium, in particular second-order $\chi^{(2)}$ and third-order $\chi^{(3)}$ nonlinearities, leading to a variety of phenomena, from optical bistability to parametric amplification and oscillation [12, 14, 13, 27, 15, 16, 28]. In the last decades, growing attention has been paid to systems featuring a complex interplay of several nonlinear phenomena at once and/or well-defined discrete modes [29, 30, 31]. Nowadays, an active frontier is to explore fundamental aspects of non-equilibrium fluctuations around the steady-state of spatially extended systems and shine a light on their relation to fundamental questions of pattern formation in nonlinear systems [32, 33, 34, 35].
@@
In this paper, we investigate a simple model based on two Kerr/Duffing oscillators coupled through nonlinear interaction terms, proportional to the product of the squares of the two amplitudes of the two modes $\sim q_{1}^{2}q_{2}^{2}$. Within the rotating wave approximation (RWA), we analyze the possible stationary paired solutions and their dynamical stability in two scenarios: one where only one of the two modes is parametrically driven, and another where both resonators are driven.
When both resonators are parametrically driven at the same frequency, the paired amplitudes of the two modes show a complex and intriguing behavior which is also characterized by the presence of several instability ranges in the frequency detuning. As a particularly interesting result, we find that when the two resonators have identical parameters and are driven with equal parametric force, the overall symmetry of the system can be spontaneously broken for moderately weak nonlinear interaction. This leads to an intriguing multistability phenomenon, featuring steady states with different amplitudes and phases on the two oscillators.
@@
The article is organized as follows. In Sec. II, we introduce the model and the effective dynamical equations in the rotating frame using the RWA. In Sec. III, we shortly discuss the behavior of the system when both resonators have non-zero oscillation amplitude when only one resonator is driven. In Sec. IV, we start analyzing the most intriguing case when both resonators are simultaneously driven and the emergence of symmetry breaking. In the successive Sec.V and Sec.VI, we respectively give an analytical form for the symmetric solution and, then, a numerical analysis for the broken symmetry case. Conclusions are finally drawn in Sec.VII.
@@ II The system
We introduce the mathematical model that describes two nonlinear parametrically driven resonantors coupled by the dispersive (cross-Kerr) interaction. We consider two resonators with equal damping coefficients and equal self Kerr-Duffing parameters. We refer to the Appendix A for further details.
In the rotating frame and using the RWA, the equations for the complex amplitudes associated with the two resonators oscillating at frequencies $\omega_{d}$ are
@@
Here $\delta\Omega_{n}$ corresponds to the scaled detuning for the resonator $n$ and $\delta\Omega_{T,n}$ is proportional to the strength of the parametric drive. It also corresponds to the (scaled) frequency threshold for the self-sustained oscillatory motion of the resonator $n$ in the limit of vanishing damping and without interaction. The parameter $g$ represents the scaled coupling strength of the nonlinear interaction.
@@
Hereafter we focus our analysis on the strong nonlinear regime in which the motion is dominated by the nonlinear interaction such that we can neglect the damping force, which is valid in the limit $3\kappa A_{n}^{2}/(8\omega_{d})\gg\Gamma$ with $A_{n}$ the amplitude of the oscillatory motion of the resonator $q_{n}(t)=A_{n}\cos(\omega_{d}t+\theta_{n})$, $\kappa$ is the Kerr-Duffiing parameter, $\Gamma$ is the damping and $\omega_{d}$ is half of the drive frequency (see appendix A). This approximation is not valid when we are close to the parametric threshold when the resonator starts to oscillate at frequency $\omega_{d}$ and the amplitude can be small enough to violate the previous condition. However, the effect of the damping is the renormalization of the threshold. Even in the presence of a nonlinear interaction, we have numerically benchmarked that small damping simply results in a minor renormalization or shift of the critical thresholds when new solutions emerge or existing solutions vanish with varying detuning.
@@
First of all, for a single parametrically driven nonlinear resonator (uncoupled case $g=0$), one has the following (stable) steady-state solution for $\bar{z}$: the resonator starts to oscillate at frequency $\omega_{d}$ above the detuning threshold $\delta\Omega>-\delta\Omega_{T,1}$ and it is entirely out of phase with respect to the drive, in the limit of vanishing damping (see appendix B). The trivial solution $\bar{z}^{(0)}_{1}=0$ is unstable in the detuning range $|\delta\Omega|<\delta\Omega_{T,1}$.
When we switch on the interaction, we discuss the stationary pair solutions. In particular, the trivial solution $\left(\bar{z}_{1}=0,\bar{z}_{2}=0\right)$ is still a stationary solution and unstable for $|\delta\Omega_{i}|<\delta\Omega_{T,i}$ as for the case $g=0$. On the other hand, for the pair solutions with the trivial solution for one of the two resonators $\bar{z}_{n}=0$, the equations for $\delta z_{1}$ and $\delta z_{2}$ become decoupled (see Eq. (4), the coefficients $c$ and $d$ vanish) but they are still correlated as the finite value of the amplitude of the driven mode $\bar{z}_{n}$ affects the fluctuation of the second non-driven mode $\delta z_{m}$ and therefore the range of stability can be different from the noninteracting case, see appendix C.
@@ III Finite amplitude solutions for one driven resonator
We now analyze the system when the external drive is applied only to the first resonator $(\delta\Omega_{T,2}=0)$. In appendix D, we give a detailed analysis of this regime, in which we report the analytical solutions in table 1. The different possible solutions are associated with different ranges of the coupling strength $g$, in particular in the ranges $30<g<1/$ 3, $11/3<g<$ 1 and $g>1$. We checked that the numerical results are in good agreement with the analytical solutions shown in appendix D.
In short, when only the first mode is driven, the cross-Kerr interaction changes the system’s stability range even when we have the trivial solution for the second non-driven mode. Furthermore, the first resonator can drive the state of the second resonator in an oscillatory motion.
We give an example of results for this regime in Fig. 1 (numerical results) in which we plot stable pair solutions (solid lines) and unstable ones (dashed lines). In particular, we notice that, for the stable solutions, the second non-driven resonator starts to oscillate for a given threshold detuning. This leads to a discontinuity in the amplitude curve of the driven resonator. Changing the parameters of the system, we also find that the curve for the amplitude of the second resonator can have a discontinuity in which the amplitude $\bar{z}_{2}$ has a jump from zero to a finite value.
Figure 1: Example of the behavior for a single-driven mode. Legend: The amplitudes of the two different resonators are distinguished by the symbol, i.e., $\blacksquare$ for $|\bar{z}_{1}|$, $\blacklozenge$ for $|\bar{z}_{2}|$. Lines with the same color show solution pairs of the two resonators. Solid lines correspond to stable solutions. Dashed lines correspond to unstable solutions. Parameters: $g=1/2$, $\delta\Omega_{T,1}=5$ and $\delta\Omega_{T,2}=0$. The blue lines show the solution pairs with only one oscillating resonator while the other one is at rest.
@@ IV.1 Numerical analysis for asymmetric drive $\delta\Omega_{T,1}\neq\delta\Omega_{T,2}$
Figure 2: Example of the behavior of the system for two driven modes with different drive strengths. Legend: The amplitudes of the two different resonators are distinguished by the symbol, i.e. $\blacksquare$ for $|\bar{z}_{1}|$, $\blacklozenge$ for $|\bar{z}_{2}|$. Lines with the same color show solution pairs of the two resonators. Solid lines correspond to stable solutions. Dashed lines correspond to unstable solutions. The green and blue lines are the paired solutions with $\bar{z}_{1}=0$ or $\bar{z}_{2}=0$. Parameters: (a) $\delta\Omega_{T,1}=3$, $\delta\Omega_{T,2}=5$, $g=1/8$. (b) $\delta\Omega_{T,1}=10/3$, $\delta\Omega_{T,2}=5$, $g=1/2$.
On the other hand, for $\delta\Omega_{T,1}\gg\delta\Omega_{T,2}$, for moderate force acting on the first resonator $\delta\Omega_{T,1}\gg 1$ and not so small coupling $g$, the solutions are connected qualitatively to ones obtained for the case of a single-driven mode, see Sec. III. One can estimate this regime in the small damping limit in which we have ${|z_{1}|}^{2}\approx\delta\Omega_{T,1}$ at small detuning. Hence the effects of the first resonator dominate on the second drive for $g\delta\Omega_{T,1}\gg\delta\Omega_{T,2}$. In such regime the interaction between the two resonators dominates over the second drive, the amplitude of the second - weakly driven - resonator is determined by its parametric interaction with the first resonator in a way qualitatively similar to the results discussed in Sec. III.
@@
For $g=1/8$ , we have stable solutions of finite amplitudes, $\bar{z}_{1}\neq 0$ and $\bar{z}_{2}\neq 0$, (turquoise lines) up to some detuning, after that they disappear, and two new stable solutions emerge again at larger detuning (purple lines). The system has multistability when we regard the phase since the modulus of the solutions shown in Fig. 2 corresponds to solutions of different phases. One can show that, in total, there are four different paired solutions corresponding to the four possible combinations of the phases.
@@
IV.2 Numerical analysis for symmetric drive $\delta\Omega_{T,1}=\delta\Omega_{T,2}$
@@
Figure 3: Example of the behavior of the system for two driven modes with symmetric drive. The orange indicates that two solutions overlap/exist with the same amplitude. The solid lines are the stable solutions, whereas the dashed lines correspond to the unstable ones. Parameters: $\delta\Omega_{T,1}=\delta\Omega_{T,2}=10$; (a) $g=1/5$; (b) $g=3/10$; (c) $g=1/2$.
@@
V Symmetric solutions for symmetric drive
$z_{0}=\pm i\,\frac{\delta\Omega+\delta\Omega_{T}}{1+3g}$ (11)
$a$ $=2\left(1+g\right){\left|z_{0}\right|}^{2}-\delta\Omega\,,$ (12)
$b$ $=-\left(1+g\right){\left|z_{0}\right|}^{2}+\delta\Omega_{T}\,,$ (13)
$c$ =±$=\pm 4g{\left|z_{0}\right|}^{2}\,,$ (14)
$d$ $=\mp 2g{\left|z_{0}\right|}^{2}\,.$ (15)
with the sign $\pm$ ()minus-or-plus($\mp$)( ∓ ) depending on the phases of the paired solution, e.g. the same $\bar{z}_{1}=\bar{z}_{2}$ or different phase $\bar{z}_{1}=-\bar{z}_{2}$. Setting $\delta z_{n}=\delta Q_{n}-i\delta P_{n}$ and changing the variables $\delta Q_{\Sigma}=\delta Q_{1}+\delta Q_{2}$, $\delta Q_{\Delta}=\delta Q_{1}-\delta Q_{2}$, $\delta P_{\Sigma}=\delta P_{1}+\delta P_{2}$, $\delta P_{\Delta}=\delta P_{1}-\delta P_{2}$, one can obtain the expansion of the Hamiltonian around the symmetric stable solution as the same of two quadratic Hamiltonian
$H_{\mathrm{RWA}}^{(2)}=H_{\Sigma}+H_{\Delta}\,.$ (16)
with
$H_{\Sigma}$ $=\frac{1}{2}\Omega_{P,\Sigma}\,\,\delta P_{\Sigma}^{2}+\frac{1}{2}\Omega_{Q,\Sigma}\,\,\delta Q_{\Sigma}^{2}\,,$ (17)
$H_{\Delta}$ $=\frac{1}{2}\Omega_{P,\Delta}\,\,\delta P_{\Delta}^{2}+\frac{1}{2}\Omega_{Q,\Delta}\,\,\delta Q_{\Delta}^{2}\,.$ (18)
in which the conjugate variables are $\left(\delta Q_{\Sigma},\delta P_{\Sigma}\right)$ and $\left(\delta Q_{\Delta},\delta P_{\Delta}\right)$ with the following dynamical equations
$\delta\dot{P}_{\Sigma}=\partial H_{\Sigma}/\partial\delta Q_{\Sigma}\,,\quad\delta\dot{Q}_{\Sigma}=-\partial H_{\Sigma}/\partial\delta\dot{Q}_{\Sigma}\,,$ (19)
$\delta\dot{P}_{\Delta}=\partial H_{\Delta}/\partial\delta\dot{Q}_{\Delta}\,,\quad\delta\dot{Q}_{\Delta}=-\partial H_{\Delta}/\delta\dot{P}_{\Delta}\,.$ (20)
The system is dynamically stable as long as the product of the two coefficients in each quadratic Hamiltonian is positive, namely $\Omega_{P,\Sigma}\Omega_{Q,\Sigma}>0$ and $\Omega_{P,\Delta}\Omega_{Q,\Delta}>0$.
For example, for $\bar{z}_{1}=\bar{z}_{2}$, the coefficients are
$\Omega_{P,\Sigma}$ $=2\left(\delta\Omega+\delta\Omega_{T}\right)\,,$ (21)
$\Omega_{Q,\Sigma}$ $=2\delta\Omega_{T}\,,$ (22)
and
$\Omega_{P,\Delta}$ $=2\,\frac{1-3g}{1+3g}\left(\delta\Omega+\delta\Omega_{T}\right)\,,$ (23)
$\Omega_{Q,\Delta}$ $=2\,\frac{\delta\Omega_{T}}{1+3g}\left[1-g\left(2\frac{\delta\Omega}{\delta\Omega_{T}}-1\right)\right]\,.$ (24)
The quadratic Hamiltonian around the stationary point $\bar{z}_{1}=-\bar{z}_{2}$ has a similar form which can be written by exchanging the coefficient $\Omega_{P,\Sigma}$ with $\Omega_{P,\Delta}$ and the coefficient $\Omega_{Q,\Sigma}$ with $\Omega_{Q,\Delta}$.
VI Broken-symmetry solutions for symmetric drive